*** I HAVE NO IDEA! ANY HELP IS APRECIATED*** Cane Company manufactures two products called Alpha and Beta that sell for $150 and $105, respectively. Each product uses only one type of raw material that costs $5 per pound. The company has the capacity to annually produce 107,000 units of each product. Its unit costs for each product at this level of activity are given below:

Alpha

Beta

  Direct materials

$

30

$

10

  Direct labor

25

20

  Variable manufacturing overhead

12

10

  Traceable fixed manufacturing overhead

21

23

  Variable selling expenses

17

13

  Common fixed expenses

20

15

  Total cost per unit

$

125

$

91

The company considers its traceable fixed manufacturing overhead to be avoidable, whereas its common fixed expenses are deemed unavoidable and have been allocated to products based on sales dollars.

13.

Assume that Cane’s customers would buy a maximum of 85,000 units of Alpha and 65,000 units of Beta. Also assume that the company’s raw material available for production is limited to 166,000 pounds. How many units of each product should Cane produce to maximize its profits?

14.

Assume that Cane’s customers would buy a maximum of 85,000 units of Alpha and 65,000 units of Beta. Also assume that the company’s raw material available for production is limited to 166,000 pounds. What is the maximum contribution margin Cane Company can earn given the limited quantity of raw materials?
Answer #1

13.

Alpha Beta
Selling price per unit $ 150 $ 105
Variable costs
Direct materials 30 10
Direct labor 25 20
Variable manufacturing overhead 12 10
Variable selling expenses 17 13
Total variable cost per unit $ 84 $ 53
Contribution margin per unit $ 66 $ 52
Pounds of raw materials per unit 6 2
Contribution margin per pound of raw material $ 11 $ 26

As the contribution margin per pound of raw material is higher for Beta, Beta should be emphasized over Alpha in production.

Maximum demand of Beta is 65,000. Raw materials required for 65,000 units of beta = 65,000 x 2 = 130,000 pounds.

Since the availability of raw materials is limited to 166,000 pounds, the remaining quantity of ( 166,000 - 130,000) = 36,000 pounds can be utilized for manufacturing Alpha. The output of Alpha would be 36,000 / 6 = 6,000 units.

Therefore, in order to maximize profits, Cane should produce 65,000 units of Beta and 6,000 units of Alpha

14. The maximum contribution that Cane can earn given the limited quantity of raw materials = 130,000 x $ 26 + 36,000 x $ 11 = $ 3,776,000